Question 1 (24 marks)
Mr Ekandjo, a law student had graduated five years ago. After graduation he decided to set up a law firm together with his close friends named: Ekandjo and Associates.
The following information apply to Ekandjo and Associates (millions of dollars) for the year 2019.
Details N$ (millions) Cash and Marketable Securities 120 Non-Current Assets 300 Sales 1050 Net Income 55
Other information
Quick ratio 2.0
Current ratio 3.0
DSO 41 days
ROE 12%
Ekandjo has no preference shares – only ordinary equity, current liabilities and long term debt.
Required:
As a Financial Management student, assist Mr Ekandjo to calculate and find the following:
a) Accounts receivable. (3 marks) b) Current liabilities (4 marks) c) Current assets (3 marks) d) Total Assets (2 marks) e) Return on Assets (ROA) (4 marks) f) Return on Equity (ROE) (4 marks) g) Long-term debt (4 marks)

Financial management
Answer #1

a ) Accounts Receivable :

1) Debtors turnover ratio = Credit sales / Average accounts receivable

    2) Debt collection period = 365 / Debtors turnover ratio

GIVEN : Debt collection period = 41 days .

Net sales = 1050 million $

Substituting the above in formula 2) ,

debtors turnover ratio = 365 / 41

= 8.9

Substituting 8.9 in formula 1) ,

   average accounts receivable = 1050 / 8.9

    ACCOUNTS RECEIVABLE = 117.98 million $

b) Current liabilities :

Quick ratio = liquid assets / current liabilities

GIVEN : Cash = 120 million $ ( liquid asset )

   Quick ratio = 2.0

Substituting the given in the above formula ,

current liabilities = 120 / 2

= 60

CURRENT LIABILITIES = 60 million $

c) Current assets :

   Current ratio = current assets / current liabilities

GIVEN : Current ratio = 3.0

current liabilities from b) = 60 million $

Substituting the given in above formula ,

Current assets = 60* 3

= 180

   CURRENT ASSETS = 180 million $

d) Total assets :

Total assets = current assets + non current assets

   GIVEN : non current assets = 300 million $

current assets from c) = 180 million $

= 180 + 300

= 480

TOTAL ASSETS = 480 million $

e) Return on assets :

   Return on assets = net income / total assets

GIVEN : net income = 55 million $

total assets from d) above = 480 million $

= 480 / 55

= 11.46%